OpenStudy (anonymous):

does any one know how to solve this problem for graphing in algebra in worded problem solving Suppose that in one week, your company takes in $2,220 in sales at one of its cell phone stores. The price breakdown per phone is as follows: Cell phone model A4: $35 Smartphone Z20: $50 Suppose that a total of 51 phones were sold.

OpenStudy (amistre64):

its a system of equations if i read it right

OpenStudy (anonymous):

yes that is correct

OpenStudy (amistre64):

you wanna know the number of phones sold?

OpenStudy (anonymous):

yes in a graph

OpenStudy (amistre64):

well graph: x + y = 51 graph: 35x +50y = 2220 and see where they cross perhaps?

OpenStudy (anonymous):

I tryed that

OpenStudy (anonymous):

and they did not cross the graph said it was x only

OpenStudy (amistre64):

#A4 = (2220 - 51(50))/(35-50) = 22

OpenStudy (amistre64):

it graphs out right; 22 and 29

OpenStudy (anonymous):

humm when I graphed it showed -0 and 2220 on the x axis

OpenStudy (amistre64):

chk yout techniques :)

OpenStudy (amistre64):

yout ..your

OpenStudy (anonymous):

I an supposed to find the x and y on the graph and don't know how

OpenStudy (anonymous):

and I did check my math

OpenStudy (amistre64):

when you graph the 2 equations; they cross at a common point; look at the site i posted for the picture. if your using a ti83 or such; go to zoomfit option